# Multiple ergodic averages along functions from a Hardy field: convergence, recurrence and combinatorial applications

### v2

@@ 5.4.   Removing the rounding function
is uniformly distributed with respect to the Haar measure $\mu_{Y_{r}}$ of $Y_{r}$ and $W$-averages. Thus, if we set $\nu\coloneqq(\mu_{Y_{0}}+\ldots+\mu_{Y_{q-1}})/q$ then it follows from 5.5 that the sequence $\big{(}\big{(}b^{f_{1}(n)},\dots,b^{f_{k}(n)}\big{)}\Gamma^{k}\big{)}_{n\in\mathbb{N}}$ is uniformly distributed with respect to $\nu$ and $W$-averages. This proves the first part of 5.7.
Next we prove that if condition 2 of A holds, then $1_{X}^{k}\in{\text{supp}}\,\nu$. Let $q_{1},\dots,q_{\ell}\in\mathbb{Z}[x]$ be jointly intersective polynomials such that $\mathrm{poly}(f_{1},\dots,f_{k})\subset\mathrm{span}^{*}(q_{1},\dots,q_{\ell})$. We claim that there exists $r\in\{0,\dots,q-1\}$ such that the polynomials $\tilde{q}_{i}:n\mapsto q_{i}(qn+r)$, $i\in\{1,\dots,\ell\}$, are jointly intersective. Indeed, since $q_{1},\dots,q_{\ell}$ are jointly intersective, for any $w\in\mathbb{N}$ there exists $n_{w}\in\mathbb{N}$ such that $q_{i}(n_{w})\equiv 0\bmod w!$ for all $i\in\{1,\dots,\ell\}$. By the pigeonhole principle, there exists $r\in\{0,\dots,q-1\}$ such that $n_{w}\equiv r\bmod q$ for infinitely many $w$. With this choice of $r$, it follows that the polynomials $\tilde{q}_{1},\ldots,\tilde{q}_{\ell}$ have a common zero modulo $w!$ for infinitely many $w\in\mathbb{N}$. But this implies that $\tilde{q}_{1},\ldots,\tilde{q}_{\ell}$ have a common zero modulo any number, proving that they are jointly intersective. We shall show that $1_{X^{k}}\in Y_{r}$.
Define, for all $i\in\{1,\ldots,k\}$, the function $\tilde{f}_{i}(t)\coloneqq f_{i}(qt+r)$, and let $\mathcal{P}\,\coloneqq\mathrm{poly}(\tilde{f}_{1},\dots,\tilde{f}_{k})$. Observe that $\mathcal{P}\subset\mathrm{span}^{*}(\tilde{q}_{1},\dots,\tilde{q}_{\ell})$. Applying 2.4 to $\tilde{f}_{1},\ldots,\tilde{f}_{k}$ we find disjoint $\mathcal{I},\mathcal{J}\subset\{1,\ldots,k\}$ with $\mathcal{I}\cup\mathcal{J}=\{1,\ldots,k\}$, $\{\lambda_{i,j}:i\in\mathcal{I},j\in\mathcal{J}\}\subset\mathbb{R}$, and $\{p_{i}:i\in\mathcal{I}\}\subset\mathcal{P}$ such that conditions 1 and 2 of 2.4 are satisfied. After reordering $\tilde{f}_{1},\ldots,\tilde{f}_{k}$ if necessary, we can assume that $\mathcal{J}=\{1,\ldots,l\}$ and =$\mathcal{I}=\{l+1,\ldots,k\}$ for some $l\in\{0,1,\ldots,k\}$ (where $l=0$ corresponds to the case $\mathcal{J}=\emptyset$ and $l=k$ to the case $\mathcal{I}=\emptyset$). For $i\in\{l+1,\ldots,k\}$ define
@@
as $n\to\infty$. Since the sequence $((b^{\tilde{f}_{1}(n)},\dots,b^{\tilde{f}_{k}(n)})\Gamma^{k})_{n\in\mathbb{N}}$ is uniformly distributed with respect to $\mu_{Y_{r}}$ and $W$-averages, it follows from (5.8) that the same is true for the sequence $\Big{(}\big{(}b^{\tilde{f}_{1}(n)},\ldots,b^{\tilde{f}_{l}(n)},b^{f_{l+1}^{*}(n)},\dots,b^{f_{k}^{*}(n)}\big{)}\Gamma^{k}\Big{)}_{n\in\mathbb{N}}$. In view of 5.11, we therefore must have that $Y_{r}=\overline{\{y_{t,n}:t\in\mathbb{R}^{l},n\in\mathbb{N}\}}$, where for $t=(t_{1},\dots,t_{l})\in\mathbb{R}^{l}$ and $n\in\mathbb{N}$ we define
@@
Finally, we assume that condition 1 of A holds. We will show that in this case $Y_{0}=X^{k}$ (and in particular $q=1$). Following the same procedure as above but with $r=0$, we conclude that $Y_{0}=\overline{\{y_{t,n}:t\in\mathbb{R}^{l},n\in\mathbb{N}\}}$ where $y_{t,n}$ is given by (5.9).
We claim that for every fixed $t\in\mathbb{R}^{l}$, the sequence $(y_{t,n})_{n\in\mathbb{N}}$ is dense in the set $\{b^{t_{1}}\Gamma\}\times\cdots\times\{b^{t_{l}}\Gamma\}\times X^{k-l}$. Indeed, this follows from [Ric23, Corollary 1.7] together with condition 1 of A (which implies that any integer linear combination $p$ of the polynomials $p_{l+1},\dots,p_{k}$ must satisfy $p-p(0)\notin\mathbb{Q}[x]$) and our assumption that $\{b^{t}\Gamma:t\in\mathbb{R}\}$ is dense in $X$. Therefore it follows that the set $\{y_{t,n}:t\in\mathbb{R}^{l},n\in\mathbb{N}\}$ is dense in $X^{k}$ and hence that $Y_{0}=X^{k}$ as desired. ∎
@@ 6.   Open questions
On the other hand, if all the $c_{i}$ are non-integers, then it follows from 1.13 that for any ergodic system $(X,\mathcal{B},\mu,T)$ and any $h_{1},\dots,h_{k}\in L^{\infty}(X)$,
@@
It may well be the case that 6.1 is true for any functions $f_{1},\ldots,f_{k}$ in a Hardy field $\mathcal{H}$.
@@
Our notion of compatibility between a tuple of functions $f_{1},\dots,f_{k}\in\mathcal{H}$ and a weight function $W$ hinges on the Property 1. This relationship is necessary to prove B because of the reliance on the results from [Ric23].  However, it is possible that a weaker notion of compatibility is sufficient. Given a Hardy field $\mathcal{H}$ and $f_{1},\dots,f_{k},W\in\mathcal{H}$ we define the property
  1. [label=Property (WP):,ref=(WP),leftmargin=*]
  2. 1.
@@
B holds if Property 1 gets replaced with Property 1.
We remark that 6.4 implies 1.12.
Condition 2 in A is somewhat complicated. In a way this is inevitable if we want to allow families of jointly intersective polynomials. However, there is a simpler, more natural, and slightly weaker condition which might be sufficient to imply the conclusion.
@@
Can one replace condition 2 in A with the weaker assumption that the collection of polynomials
[…]
@@
One of the motivations for this paper was to expand on our previous work from [BMR20] which revealed a new phenomenon pertaining to non-polynomial functions, say, from a Hardy field (see 1.6 above). As a corollary of 1.6, if $f\in\mathcal{H}$ and $t^{\ell-1}\prec f(t)\prec t^{\ell}$ for some $\ell\in\mathbb{N}$, then for every $E\subset\mathbb{N}$ with $\bar{d}(E)>0$ the set
@@
$R_{f}\text{ is }\begin{cases}\text{ thick }&\text{ if }t^{\ell-1}\prec f(t)\prec t^{\ell}\text{ for some }\ell\in\mathbb{N}\\ \text{ syndetic }&\text{ otherwise}.\end{cases}$
Corollary A4 in the Introduction implies that if $f_{1},\dots,f_{k}$ belong to a Hardy field and satisfy a “non-polynomiality condition” then the intersection $R_{f_{1}}\cap\cdots\cap R_{f_{k}}$ is thick. While we can not replace the condition with the more natural $\mathrm{poly}(f_{1},\dots,f_{k})=\emptyset$, as shown in 1.19, there is an intermediate condition which might be sufficient.
@@
[…]

### v3

@@ 5.4.  Removing the rounding function
is uniformly distributed with respect to the Haar measure $\mu_{Y_{r}}$ of $Y_{r}$ and $W$-averages. Thus, if we set $\nu\coloneqq(\mu_{Y_{0}}+\ldots+\mu_{Y_{q-1}})/q$ then it follows from ˜5.5 that the sequence $\big(\big(b^{f_{1}(n)},\dots,b^{f_{k}(n)}\big)\Gamma^{k}\big)_{n\in\mathbb{N}}$ is uniformly distributed with respect to $\nu$ and $W$-averages. This proves the first part of ˜5.7.
Next we prove that if condition (2) of Theorem A holds, then $1_{X}^{k}\in{\text{supp}}\,\nu$. Let $q_{1},\dots,q_{\ell}\in\mathbb{Z}[x]$ be jointly intersective polynomials such that $\mathrm{poly}(f_{1},\dots,f_{k})\subset\mathrm{span}^{*}(q_{1},\dots,q_{\ell})$. We claim that there exists $r\in\{0,\dots,q-1\}$ such that the polynomials $\tilde{q}_{i}:n\mapsto q_{i}(qn+r)$, $i\in\{1,\dots,\ell\}$, are jointly intersective. Indeed, since $q_{1},\dots,q_{\ell}$ are jointly intersective, for any $w\in\mathbb{N}$ there exists $n_{w}\in\mathbb{N}$ such that $q_{i}(n_{w})\equiv 0\bmod w!$ for all $i\in\{1,\dots,\ell\}$. By the pigeonhole principle, there exists $r\in\{0,\dots,q-1\}$ such that $n_{w}\equiv r\bmod q$ for infinitely many $w$. With this choice of $r$, it follows that the polynomials $\tilde{q}_{1},\ldots,\tilde{q}_{\ell}$ have a common zero modulo $w!$ for infinitely many $w\in\mathbb{N}$. But this implies that $\tilde{q}_{1},\ldots,\tilde{q}_{\ell}$ have a common zero modulo any number, proving that they are jointly intersective. We shall show that $1_{X^{k}}\in Y_{r}$.
[…]
@@
as $n\to\infty$. Since the sequence $((b^{\tilde{f}_{1}(n)},\dots,b^{\tilde{f}_{k}(n)})\Gamma^{k})_{n\in\mathbb{N}}$ is uniformly distributed with respect to $\mu_{Y_{r}}$ and $W$-averages, it follows from (5.8) that the same is true for the sequence $\Big(\big(b^{\tilde{f}_{1}(n)},\ldots,b^{\tilde{f}_{l}(n)},b^{f_{l+1}^{*}(n)},\dots,b^{f_{k}^{*}(n)}\big)\Gamma^{k}\Big)_{n\in\mathbb{N}}$. In view of ˜5.11, we therefore must have that $Y_{r}=\overline{\{y_{t,n}:t\in\mathbb{R}^{l},n\in\mathbb{N}\}}$, where for $t=(t_{1},\dots,t_{l})\in\mathbb{R}^{l}$ and $n\in\mathbb{N}$ we define
@@
Finally, we assume that condition (1) of Theorem A holds. We will show that in this case $Y_{0}=X^{k}$ (and in particular $q=1$). Following the same procedure as above but with $r=0$, we conclude that $Y_{0}=\overline{\{y_{t,n}:t\in\mathbb{R}^{l},n\in\mathbb{N}\}}$ where $y_{t,n}$ is given by (5.9).
We claim that for every fixed $t\in\mathbb{R}^{l}$, the sequence $(y_{t,n})_{n\in\mathbb{N}}$ is dense in the set $\{b^{t_{1}}\Gamma\}\times\cdots\times\{b^{t_{l}}\Gamma\}\times X^{k-l}$. Indeed, this follows from [Ric23, Corollary 1.7] together with condition (1) of Theorem A (which implies that any integer linear combination $p$ of the polynomials $p_{l+1},\dots,p_{k}$ must satisfy $p-p(0)\notin\mathbb{Q}[x]$) and our assumption that $\{b^{t}\Gamma:t\in\mathbb{R}\}$ is dense in $X$. Therefore it follows that the set $\{y_{t,n}:t\in\mathbb{R}^{l},n\in\mathbb{N}\}$ is dense in $X^{k}$ and hence that $Y_{0}=X^{k}$ as desired. ∎
@@ 6.  Open questions
On the other hand, if all the $c_{i}$ are non-integers, then it follows from ˜1.13 that for any ergodic system $(X,\mathcal{B},\mu,T)$ and any $h_{1},\dots,h_{k}\in L^{\infty}(X)$,
@@
It may well be the case that ˜6.1 is true for any functions $f_{1},\ldots,f_{k}$ in a Hardy field $\mathcal{H}$.
@@
Our notion of compatibility between a tuple of functions $f_{1},\dots,f_{k}\in\mathcal{H}$ and a weight function $W$ hinges on the Property (P). This relationship is necessary to prove Theorem B because of the reliance on the results from [Ric23]. However, it is possible that a weaker notion of compatibility is sufficient. Given a Hardy field $\mathcal{H}$ and $f_{1},\dots,f_{k},W\in\mathcal{H}$ we define the property
  1. Property (WP):
@@
Theorem B holds if Property (P) gets replaced with Property (WP).
We remark that ˜6.4 implies ˜1.12.
Condition (2) in Theorem A is somewhat complicated. In a way this is inevitable if we want to allow families of jointly intersective polynomials. However, there is a simpler, more natural, and slightly weaker condition which might be sufficient to imply the conclusion.
@@
Can one replace condition (2) in Theorem A with the weaker assumption that the collection of polynomials
$\mathrm{poly}(f_{1},\ldots,f_{k})\cap\mathbb{Z}[t]\,=\,\Big\{q\in\mathbb{Z}[t]:\exists f\in\mathrm{span}^{*}(f_{1},\dots,f_{k})\,\text{with}\,\lim_{t\to\infty}|f(t)-q(t)|=0\Big\}$
@@
One of the motivations for this paper was to expand on our previous work from [BMR20] which revealed a new phenomenon pertaining to non-polynomial functions, say, from a Hardy field (see ˜1.6 above). As a corollary of ˜1.6, if $f\in\mathcal{H}$ and $t^{\ell-1}\prec f(t)\prec t^{\ell}$ for some $\ell\in\mathbb{N}$, then for every $E\subset\mathbb{N}$ with $\bar{d}(E)>0$ the set
@@
[…]
Corollary A4 in the Introduction implies that if $f_{1},\dots,f_{k}$ belong to a Hardy field and satisfy a “non-polynomiality condition” then the intersection $R_{f_{1}}\cap\cdots\cap R_{f_{k}}$ is thick. While we can not replace the condition with the more natural $\mathrm{poly}(f_{1},\dots,f_{k})=\emptyset$, as shown in ˜1.19, there is an intermediate condition which might be sufficient.
@@
In particular, we don’t know if letting $f_{1}(t)=t^{3/2}$ and $f_{2}(t)=\alpha t^{3/2}+t$, where $\alpha\in\mathbb{R}\backslash\mathbb{Q}$, the intersection $R_{f_{1}}\cap R_{f_{2}}$ is thick.